### v2

# Photoinduced Charge Transfer in Transition Metal Dichalcogenide Quantum Dots

Abstract
In this paper, we have explored the charge transfer mechanism in transition metal dichalcogenide (TMDC) quantum dots (QDs) of molybdenum disulfide () and tungsten disulfide (). Rhodamine 6G (R6G), a dye from the rhodamine family, has been employed as the fluorescent molecule, with MoS2 and WS2 QDs acting as electron acceptors in the photo-induced charge transfer process. The TMDC QDs were synthesized using a top-down approach and characterized through transmission electron microscopy (TEM), UV-Vis spectrophotometry, and fluorimetry. TEM images revealed well-dispersed particles measuring 2 nm in size. These QDs exhibit strong fluorescence emission when excited with light at wavelengths below 350 nm. Under light exposure, photons generate charges in the fluorescent dye molecules, and the TMDC QDs facilitate the charge transfer process. The charge transfer phenomenon was investigated using time-correlated single photon counting (TCSPC), a time-resolved fluorescence spectroscopic technique. The time-resolved fluorescence study indicated a change in the fluorescence (FL) lifetime of R6G molecules in the presence of QDs. The FL lifetime of R6G molecules without QDs was found to be 4.0 ns, which decreased to 1.9 ns and 3.8 ns in the presence of MoS2 and WS2 QDs, respectively. This reduction in FL lifetime suggests that the MoS2 and WS2 QDs provide an additional pathway for photo-generated electrons in the excited state of R6G molecules. This research can be extended to optoelectronic devices, where charge transfer is crucial for device efficiency and performance.
In recent years, different kinds of quantum dots (QDs) have been extensively researched because of their numerous applications in cell imaging, photovoltaic devices, light-emitting devices, and more. Xu et al. (2021); Li et al. (2024); Wang et al. (2020); Sadeghi et al. (2019); Li et al. (2016); Xue et al. (2019); Meng et al. (2024) The most appealing characteristic of these materials is their size-dependent opto-electronic properties, making them ideal candidates for designing various opto-electronic devices.Li et al. (2015); Litvin et al. (2017) The fascinating properties of QDs have been investigated in composite systems or devices through their combination with various materials, including molecules, other QDs, and 2D nanomaterials. When QDs are paired with biomolecules, they are utilized for bio-imaging, while QDs in a core-shell configuration are explored for optoelectronic devices. These QDs, when sensitized with other semiconductors, have been used to create cost-effective, next-generation photovoltaic devices for various applications. QDs are favored in these applications due to their size-dependent optical and electronic properties. The interaction between QDs and their surrounding media or materials is critical for device performance. Although the electronic interactions between organic molecules and fluorescent QDs have been well studied, systems where QDs are coupled with other inorganic molecules have been less explored. These interactions differ from those in QD-molecular systems because inorganic materials have a continuum of electronic states, whereas molecular acceptors have discrete electronic states. When QDs are coupled with other inorganic materials, they can donate electrons, as observed in photovoltaic devices, or act as electron donors/acceptors, as seen in light-emitting diodes (LEDs). In these applications, electron transfer involves a specific direction of transfer. Therefore, understanding the factors driving electron transfer in these composite systems is crucial for better comprehending and further exploiting the unique properties of QDs in optoelectronic devices.Algar and Krull (2008); Wang et al. (2021); Chakraborti et al. (2016); Zhu, Zhang, and Xia (2018); Lima et al. (2022); Wu et al. (2020); Rani et al. (2020)
There are various studies on different kind of QDs, where these QDs have been used as acceptor/donor or both of the electrons. They have also been explored for donor or acceptor in the resonance energy transfer. In recent years researchers have explored various kinds of layered based QDs in place of traditional QDs, e.g., CdS, CdSe, PbS, etc.Kharangarh et al. (2023); Sariga et al. (2023); Gan et al. (2015); Guo et al. (2019) In optoelectronic applications, QDs function as light-harvesting and charge-separation centers. A crucial step in this process is the transfer of electrons or holes from the photoexcited QDs to nearby acceptors through a nanoscale interface. The overall efficiency of the device relies on these charge transfer processes, which are influenced by competing loss pathways. Therefore, understanding and controlling interfacial charge transfer processes is vital for improving QD-based devices. The size of the QDs also significantly impacts the charge transfer process, with studies showing that the electron transfer rate increases as the QD size decreases. Zhu et al. (2016)
In recent years, various 2D layered nanomaterials have been transformed into QDs, which exhibit significantly different properties compared to their 2D structures and bulk counterparts. Specifically, 2D layered transition metal dichalcogenides (TMDCs) have been converted into QDs using various methods. Among these TMDC-based QDs, MoS2 and WS2 QDs have been extensively studied. They exhibit intriguing optical and electronic properties, with their electronic structure being tunable by adjusting their size and surface morphology. These QDs have been applied in various fields, including sensors, LEDs, photodetectors, and cell imaging. Yan et al. (2016); Ghorai et al. (2017); Singh et al. (2019); Guo and Li (2020)
Several studies have used TMDC QDs as energy donors in hybrid systems due to their photoluminescence in the UV region. There is extensive research on energy transfer and charge transfer in various nanomaterials and their applications in sensors, photodetectors, and solar cells. In this work, we have investigated charge transfer in two different types of fluorescent materials: fluorescent dye molecules and QDs. Typically, QDs are used as donors or acceptors, and dye molecules as acceptors or donors in energy transfer studies, which often focus on the energy transfer and its applications. Here, we have explored the potential for charge transfer from dye molecules to QDs, providing valuable insights for developing optoelectronic devices based on fluorescent nanomaterials. In this paper, we used a common dye molecule, R6G, as the source of photo-excited electrons and TMDC QDs as the electron acceptors.
In order to synthesize the QDs and perform this study, various chemicals were used. Molybdenum disulphide (MoS2), tungsten disulphide (WS2) powder (2m size), N-Methyl 2-pyrrolidone (NMP), and Rhodamine 6G dye were purchased from Sigma Aldrich. All the chemicals were used as received without further purification.
The TMDC QDs were synthesized using top-down strategy as mentioned in Long et al. with slight modification.Long et al. (2016) In typical synthesis procedure, 50 mg of TMDC powder was dissolved in 50 mL of NMP solution with constant stirring for 15 hours at 29 ∘C. The resulting dark suspension was centrifuged at 10,000 rpm for 40 min. The supernatant of the solution was collected and used for further studies. The schematic diagram of the synthesis of TMDC quantum dots is shown in Fig 1.
In our study, prior to exploring charge transfer mechanism, the quantum dots (QDs) were characterized using transmission electron microscopy (TEM), fluorescence spectroscopy, and absorption spectroscopy. TEM images were captured with a JEOL-2010 TEM operating at 200 kV. The absorption spectra of the QDs were obtained using a Shimadzu (UV-2450) UV-Vis spectrophotometer. In order to study charge transfer in these TMDCs QDS, the QDs were mixed with a fixed quantity of R6G solution. For charge transfer studies, 0.5 mL aqueous solution of 10-6 M R6G was further diluted to 3.5 mL, and 1 mL of QDs solution was mixed to R6G solution. In all the samples the NMP and water contents were kept same. The steady-state fluorescence spectra of R6G molecules were captured both in the absence and presence of TMDCs QDs using a fluorimeter with an excitation wavelength of 430 nm. The time-resolved fluorescence analysis was conducted using a time-correlated single-photon counting module (TCSPC) (FL920, Edinburgh Instruments, UK), with excitation provided by a diode laser at 465 nm.
The absorption spectra of MoS2 and WS2 quantum dots are shown in Fig 2. A clear absorption peak around 325 nm is observed, which is a characteristic peak of WS2 quantum dots and it is due to excitonic features of the QDs. Similarly, for MoS2 QDs a band edge around 380 nm is observed. The color of suspension and the absorption peaks are similar to the previous reported results. Coleman et al. (2011); Long et al. (2016) When the MoS2 and WS2 are in the bulk or nanosheet form, they exhibit quite different absorption patterns having two distinct peaks around 600 nm which are known as and excitonic peaks. While in case of QDs these and excitonic peaks disappear and absorption is observed in UV region. The pronounced blue shift observed in the absorption spectra of the synthesized MoS2 and WS2 QDs can be attributed to the quantum confinement effect and edge effects that become significant as the lateral dimensions of the QDs are reduced to a few tens of nanometers. Li et al. (2017); Liu et al. (2014) In order to have insight of the band gaps of these QDs, the Tauc plots of these QDs are drawn. The Tauc plots of the quantum dots are shown in Fig 3.
In case of semiconductors, it has been observed the bandgap increases with decrease in the size of the semiconducting quantum dots. The size dependent bandgap is given by following equation:Wu et al. (2005)
In this expression, represents the bandgap of the bulk material, is Planck’s constant, is the reduced mass of the exciton, is the quantum dot radius, is the electron charge, is the permittivity of free space, and is the dielectric constant of the material. This approximation assumes the particles are spherical, and the motions of the electron and hole are described by their effective masses. The effect of surface atoms are also ignored in this approximation. For MoS2, the reduced mass of exciton is 0.27 times that of the free electron, the bulk bandgap is 1.29 eV, and the permittivity is about 6.8 times that of free space. The calculated value of the bandgap for MoS2 quantum dots is found to be 1.9 eV. Similarly for WS2 the bandgap was found to be 2.7 eV. In case of WS2, the reduced mass was taken as 0.146 times that of the free electron mass, the dielectric constant of the material was taken as 4.13 while theoretically calculation of the band gap. In the Equation 1, the second and third terms of the equation are the kinetic energy and electrostatic attraction terms, respectively. It is the kinetic energy term which dominates in general, and results a blue-shift in the bandgap as the size of the quantum dots decreases. As the size of QDs decreases, the bandgap increases because the valence band shifts to more positive potentials, while the conduction band shifts to more negative potentials. In most semiconducting QDs, this increase in bandgap is primarily due to the conduction band moving to more negative potentials, with only a slight shift in the valence band. (considering potential relative to NHE). Gan et al. (2015); Wu et al. (2005); Parsapour et al. (1996) The calculated bandgap is found to be different from the experimentally observed values using Tauc plot which is due to the different size of the QDs, hence there might be variation in the reduced mass of the excitons and the variation in the dielectric constants values at the nanometer scale.
The fluorescence spectra of MoS2, and WS2 are shown in Fig. 4. The QDs show emission in blue region of the electromagnetic spectrum. The origin of this fluorescence emission is quantum confinement effect and the diract bandgap transition in these quantum dots. The emission spectrum has a wide pattern which is due to variation in the size of these QDs. The distribution of these particles size is having broad nature. Ha et al. (2014)
The structural properties of MoS2 and WS2 quantum dots were explored by using transmission electron microscopy (TEM). Fig. 5 shows the TEM images of the quantum dots. The average size of QDs is of the order of 1.2 nm in case of WS2 QDs, and 1.7 nm for the MoS2 QDs.
The photo-response of R6G molecules, both in the absence and presence of QDs, was investigated using fluorescence spectroscopy. The steady-state fluorescence spectra of R6G molecules, with and without MoS2 and WS2 QDs, were recorded using a fluorimeter with a 430 nm excitation wavelength. Emission measurements were taken across a wavelength range of 520-620 nm. The steady state fluorescence spectra of R6G molecules in the absence and presence of MoS2, and WS2 are shown in Fig. 6. The R6G molecules show characteristic emission peak around 550 nm.
In Fig. 6, it was observed that the fluorescence intensity of R6G decreases in the presence of QDs, indicating quenching by the QDs. This quenching could be either dynamic or static, which can be confirmed through time-resolved fluorescence studies. To investigate potential charge transfer from R6G to the QDs, the fluorescence lifetime of R6G molecules was measured both with and without the QDs using a time-correlated single photon counting (TCSPC) setup. The resulting fluorescence lifetime decay spectra of R6G in both scenarios are depicted in the Fig. 7.
Fig. 7 presents the fluorescence lifetime decay curves of R6G molecules both in the presence of MoS2 and WS2 quantum dots (QDs) and in their absence. These decay curves were analyzed using exponential decay functions of the following form: Patel, Sahoo, and Mohanty (2014)
The expression for the fluorescence decay includes as the number of discrete decay components, with representing the decay times, and as the weighting factors corresponding to the decay component. It was observed that the fluorescence decay curves steepened in the presence of quantum dots (QDs) compared to pure R6G. While the R6G decay curve alone was well-fitted with a single-exponential function (), in the presence of QDs, a bi-exponential function () was required for fitting. The average fluorescence lifetime for R6G alone was 4.0 ns, but this decreased to 1.9 ns with MoS2 QDs and 3.8 ns with WS2 QDs. The reduction in fluorescence lifetime suggests charge or energy transfer from R6G molecules to the QDs, with the effect being more pronounced for MoS2 QDs. To further investigate the reasons for this decrease in fluorescence lifetime, the absorption spectra of the QDs were compared with the emission spectrum of R6G molecules, as depicted in Fig. 8.
In Fig. 8, the emission spectrum of R6G does not overlap with the absorption spectra of TMDC QDs, which rules out energy transfer from R6G to the QDs as a cause for the observed reduction in fluorescence intensity and lifetime. Instead, the decrease is likely due to charge transfer from photo-excited R6G molecules to the QDs. This charge transfer process leads to quenching of both fluorescence intensity and lifetime. The charge transfer efficiency from R6G to TMDC QDs was calculated using the following equation:
Here, and are fluorescence lifetimes of R6G molecules, in absence and presence of TMDCs QDs (acceptors), respectively. The charge transfer efficiency, in case of MoS2 QDs is found to be 53%. For WS2, the efficiency is found to be 3 %.
The mechanism of charge transfer from R6G molecules to TMDCs QDs is illustrated in Fig. 9, and Fig. 10. The schematic of energy levels of R6G, MoS2 and WS2 quantum dots are shown in Fig. 9, and a schematic of charge transfer in the QDs are shown in Fig. 10.
It has been previously observed that the dye can be brought to excited state by absorbing the photons of suitable wavelength and further the excited dye can transfer electrons to the conduction band of adjacent semiconducting materials, in this way the dye can be degraded by the reactive oxygen species.Li et al. (2010) Here, in this study, we used R6G dye molecules which has redox potentials 1.2 V and -1.1 V vs NHE for ground state (HOMO) and excited state (LUMO), respectively.Li et al. (2010) The energy levels of R6G in its ground state and excited state versus vacuum is calculated using equation E(eV) = -4.5 -ENHE and the respective energy levels are -5.7 eV and -3.4 eV, for ground and excited state, respectively. The optical band gap of R6G dye molecules is of the order of 2.3 eV, which corresponds to 540 nm of the wavelength. The potential of the valence band and conduction band of MoS2 are found to be -6.0 eV and -4.2 eV, respectively. Similarly, the valence band and conduction band energy levels of WS2 are found to be at -5.48 eV, and -3.93 eV, respectively. Mukherjee et al. (2016) From these energy levels, it is observed that by absorbing photons of wavelength 530 nm, R6G molecules make transition into an excited state R6G∗ and the energy differences between R6G* state to the conduction bands of QDs provide a thermodynamically favourable condition for the transfer of electrons from LUMO of R6G to the conduction band of MoS2 and WS2 QDs. The energy difference between LUMO of R6G and conduction band of MoS2 is 0.8 eV and for WS2, the difference is 0.53 eV. The energy differnce is maximum in case of MoS2 which favours the charge transfer from R6G molecules in the excited state to the conduction band of the MoS2 as compared to the WS2. Therefore the charge transfer efficiency is maximum (53) in MoS2 QDs. This study can be further extended to other quantum dots and 2D materials and the efficiency of charge transfer can be tailored by selecting appropriate materials.
In summary, the MoS2 and WS2 QDs were synthesized using chemical assisted exfoliation method. The photo-induced charge transfer between fluorescent R6G molecules and TMDC QDs of MoS2 and WS2 was studied using fluorescence spectroscopic techniques. The QDs of MoS2 and WS2 were used as acceptors for charge transfer, which could be further explored to developing QDs-molecules based optoelectronic devices. Maximum charge transfer efficiency was observed in case of MoS2 QDs which was due to a wide energy difference between the LUMO of the R6G molecules and the conduction energy level of the MoS2 QDs, it had provided a thermodynamically favourable condition of the charge transfer.
Authors acknowledge the AIRF, JNU for TRFS, and TEM characterizations. Authors are also thankful to Dr. Jaydeep Bhattacharya of the School of Biotechnology, JNU for absorption and fluorescence studies.